A study of iron deficiency among infants compared samples of infants following different feeding regimens. One group contained breast-fed infants, while the children in other group were fed a standard baby formula without any iron supplements. Here are summary results on blood hemoglobin levels at 12 months of age.


Group n Mean s
Breast-fed 23 13.3 1.7
Formula 19 12.4 1.8

a.) Is there significant evidence that the mean hemoglobin level is higher among breast-fed babies? State null hypothesis and alternate hypothesis and conduct at –test.
b.) Give a 95% confidence interval for the mean difference in hemoglobin level between the two populations of infants.


qus 2

A small accounting firm pays each of its five clerks Rs 25000, two junior accountants Rs 60000, and the firm‘s owner Rs 255000. What is the mean salary paid at this Firm? How many of the employees earn less than the mean? What is the median salary?



Answer by stanbon(48568)

A study of iron deficiency among infants compared samples of infants following different feeding regimens. One group contained breast-fed infants, while the children in other group were fed a standard baby formula without any iron supplements. Here are summary results on blood hemoglobin levels at 12 months of age.

Group...... n Mean. s
Breast-fed 23 13.3 1.7
Formula... 19 12.4 1.8
a.) Is there significant evidence that the mean hemoglobin level is higher among breast-fed babies? State null hypothesis and alternate hypothesis and conduct a t –test.
Ho: u(breast)-u(formula) = 0
Ha: u(breast)-u(formula) > 0
---
I ran a 2-Sample T-test and got the following results:
test statistic: t = 1.2862
p-value: 0.1031
df = 37.5976
Conclusion: Since the p-value is greater than 5%, Fail to reject Ho.
---------------------------------------------------------------------------
b.) Give a 95% confidence interval for the mean difference in hemoglobin level between the two populations of infants.
---
I ran a 2-Sample T-Test Interval program of got the following:
95% C.I.: -0.4021< u(b)-u(f) < 1.8021
=====================================================
A small accounting firm pays each of its five clerks Rs 25000, two junior accountants Rs 60000, and the firm‘s owner Rs 255000. What is the mean salary paid at this Firm? How many of the employees earn less than the mean? What is the median salary?
mean = 500000/8 = 62500
# less than mean is 7 employees
median 25000
=======================================
Cheers,
Stan H.